OpenStudy (anonymous):

What is the resistance of a lightbulb that uses an average power of 100.0 watts when connected to a 60.0 hz power source having a maximum voltage of 156 volts

OpenStudy (anonymous):

my approach is \[power = (\Delta V)^2/R\] \[R = (\Delta V)^2/Power\] \[R = (156)^2/100\] however i don't think this is right. can anyone point me in the right direction?